# Few-Shot Adversarial Low-Rank Fine-Tuning of Vision-Language Models

### v2

@@ 3.1 Adversarial Fine-Tuning of CLIP via LoRA
where $\Delta$ is a bounded set of admissible perturbations, and $f:\mathbb{R}^{d\times k+n}\rightarrow\mathbb{R}$ is a non-convex loss function expressible in the stochastic form $\mathbb{E}_{\xi\sim\mathcal{D}}[F(W_{0}+BA,\delta;\xi)]$. Here, the expectation is taken over randomly sampled batches $\xi\sim\mathcal{D}$, where $\mathcal{D}$ represents the underlying data distribution.
@@ 3.2 AdvCLIP-LoRA Algorithm
In this section, we present the proposed AdvCLIP-LoRA algorithm, which solves the minimax problem (4) to enhance the adversarial robustness of a CLIP model fine-tuned with LoRA. The AdvCLIP-LoRA algorithm proceeds for $T$ iterations. At each iteration $t$:
1) Select $M$ independent and identically distributed (i.i.d.) samples $\{\xi_{i}\}_{i=1}^{M}$ from the dataset.
2) Update the perturbation $\delta$ via:
$\delta_{t}=\mathcal{P}_{\Delta}\left(\delta_{t-1}+\frac{\eta_{\delta}}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i})\right),$ (5)
where $\eta_{\delta}$ is the learning rate for $\delta$, $\Delta$ is a bounded perturbation set, and $\mathcal{P}_{\Delta}$ projects onto $\Delta$.
3) Update the LoRA matrices $A$ and $B$ using the current $\delta_{t}$:
$A_{t}$ $=A_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{A}F(W_{t-1},\delta_{t};\xi_{i})\right),$
$B_{t}$ $=B_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{B}F(W_{t-1},\delta_{t};\xi_{i})\right),$ (6)
where $\eta_{w}$ is the learning rate for $A$ and $B$. In the end, the algorithm randomly draws $\hat{A},\hat{B}$ from $\{A_{t},B_{t}\}_{t=1}^{T}$ uniformly at random. [^1] This is a standard practice in nonconvex optimization for stochastic gradient descent to find stationary points. In practice, we select the model that achieves the highest accuracy on the validation set. The steps of the AdvCLIP-LoRA algorithm are illustrated in Fig. 2. Moreover, the AdvCLIP-LoRA pipeline can be found in Alg. 1.
@@ 4 Convergence Analysis
In this section, we present a thorough convergence analysis of the proposed AdvCLIP-LoRA algorithm. The complete proofs can be found in Appendix B. Let us begin with a few definitions.
Definition 4.1
A function $f$ is $L$-Lipschitz if for all $W,W^{\prime}$, we have
Definition 4.2
A function $f$ is $\ell$-smooth if for all $W,W^{\prime}$, we have
Consider the minimax problem (4), which is equivalent to minimizing the function $\Phi(\cdot)=\max_{\delta\in\Delta}f(\cdot,\delta)$. In the context of nonconvex-strongly-concave minimax problems, where $f(W,\cdot)$ is strongly-concave for each $W$, the maximization problem $\max_{\delta\in\Delta}f(W,\delta)$ can be solved efficiently, yielding useful insights into $\Phi$. However, finding the global minimum of $\Phi$ remains NP-hard in general due to its nonconvex nature. To address this challenge, we define local surrogates for the global minimum of $\Phi$. One commonly used surrogate in nonconvex optimization is the notion of stationarity, which is suitable when $\Phi$ is differentiable.
Definition 4.3
A point $W$ is an $\epsilon$-stationary point $(\epsilon\geq 0)$ of a differentiable function $\Phi$ if $\|\nabla\Phi(W)\|\leq\epsilon$.
@@
The objective function and constraint set $\left(f:\mathbb{R}^{d\times k+n}\rightarrow\mathbb{R},\Delta\subseteq\mathbb{R}^{n}\right)$ satisfy
  1. 1.
    $f$ is $\ell$-smooth and $f(W,\cdot)$ is $\mu$-strongly concave.
  2. 2.
    $\Delta$ is a convex and bounded set with a diameter $D\geq 0$.
Assumption 4.6
Let $W_{t}=W_{0}+B_{t}A_{t}$ represent the model parameters at the $t$-th step. For all $t=1,\cdots,T$ there exist constants $c_{A}>0$ and $c_{B}>0$ such that: $\left\|A_{t}\right\|_{F}\leq c_{A}$ and $\left\|B_{t}\right\|_{F}\leq c_{B}$.
We next present a proposition on the structure of the function $\Phi$. Let $\kappa=\ell/\mu$ denote the condition number and define
@@
Proposition 4.7
[36] Under Assumption 4.5, $\Phi(\cdot)$ is $2\kappa\ell$-smooth with $\nabla\Phi(\cdot)=\nabla_{W}f\left(\cdot,\delta^{\star}(\cdot)\right)$. Also, $\delta^{\star}(\cdot)$ is $\kappa$-Lipschitz.
Using Proposition 4.7 and Assumption 4.6, we can prove the smoothness of $\Phi(\cdot)$ with respect to $A$ and $B$ when the other is held fixed. Formally, we state the following lemma:
Lemma 4.8
Under Assumptions 4.5 and 4.6, the function $\Phi(\cdot)$ is $2\kappa\ell c_{B}^{2}$-smooth with respect to $A$ when $B$ is fixed and $2\kappa\ell c_{A}^{2}$-smooth with respect to $B$ when $A$ is fixed.
Now, we present the main theoretical results using the order-wise notation for AdvCLIP-LoRA.
@@
Under Assumptions 4.4, 4.5, and 4.6, and letting the learning rates be chosen as
$\eta_{w}=\Theta\left(\mathrm{min}\left\{\tfrac{1}{\kappa\ell(c_{A}^{4}+c_{B}^{4})},\tfrac{1}{\kappa^{2}\ell(c_{A}^{2}+c_{B}^{2})},\tfrac{1}{(G^{2}+\kappa\ell c_{A}^{4}c_{B}^{2})^{1/2}}\right\}\right),$ (11)
and $\eta_{\delta}=\Theta(1/\ell)$, the number of iterations required by AdvCLIP-LoRA to return an $\epsilon$-stationary point is bounded by
where $\Delta_{\Phi}=\mathbb{E}\Phi(W_{0})-\mathbb{E}\Phi(W_{T+1})$. Moreover, the mini-batch size $M$ is bounded by
@@
First, AdvCLIP-LoRA is guaranteed to find an $\epsilon$-stationary point of $\Phi(\cdot)$ within $\mathcal{O}\left(\epsilon^{-2}\right)$ iterations and with a total stochastic gradient complexity of $\mathcal{O}\left(\epsilon^{-4}\right)$. Moreover, our choice of learning rates satisfies $\eta_{w}\ll\eta_{\delta}$, which stems from the non-convex–strong-concave structure of the objective function.

### v3

@@ 3.1 Adversarial Fine-Tuning of CLIP via LoRA
where $\Delta$ is a bounded set of admissible perturbations, and $f:\mathbb{R}^{d\times k+n}\rightarrow\mathbb{R}$ is a non-convex loss function expressible in the stochastic form $\mathbb{E}_{\xi\sim\mathcal{D}}[F(W_{0}+BA,\delta;\xi)]$. Here, the expectation is taken over sampled batches $\xi\sim\mathcal{D}$, where $\mathcal{D}$ represents the underlying data distribution.
@@ 3.2 AdvCLIP-LoRA Algorithm
In this section, we present the proposed AdvCLIP-LoRA algorithm, which solves the minimax problem (Eq. 5) to enhance the adversarial robustness of a CLIP model fine-tuned with LoRA. The AdvCLIP-LoRA algorithm proceeds for $T$ iterations. At each iteration $t$:
1) Select $M$ samples $\{\xi_{i}\}_{i=1}^{M}$ from the dataset.
2) Update the perturbation $\delta$ for $\tau$ iterations via:
$\delta_{t}=\mathcal{P}_{\Delta}(\delta_{t-1}+\frac{\eta_{\delta}}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i})),$ (6)
where $\eta_{\delta}$ is the learning rate for $\delta$, $\Delta$ is a bounded perturbation set, and $\mathcal{P}_{\Delta}$ projects onto $\Delta$. The set $\Delta$ may be any convex, bounded subset of $\mathbb{R}^{n}$; in our experiments we take $\Delta=\{\delta:\|\delta\|_{\infty}\leq\epsilon\}$, i.e., an $\ell_{\infty}$-ball of radius $\epsilon$.
3) Update the LoRA matrices $A$ and $B$ using the current $\delta_{t}$ to obtain $A_{t}$ and $B_{t}$ (lines 8 and 9 of Alg. 1), where $\eta_{w}$ is the learning rate for AA and $B$. The steps of the AdvCLIP-LoRA algorithm are illustrated in Fig. 2. Moreover, the AdvCLIP-LoRA pipeline can be found in Alg. 1.
@@ 4 Convergence Analysis
In this section, we present a thorough convergence analysis of the proposed AdvCLIP-LoRA algorithm. The complete proofs can be found in Appendix C.
Consider the minimax problem (Eq. 5), which is equivalent to minimizing the function $\Phi(\cdot)=\max_{\delta\in\Delta}f(\cdot,\delta)$. In the context of nonconvex-strongly-concave minimax problems, where $f(W,\cdot)$ is strongly-concave for each $W$, the maximization problem $\max_{\delta\in\Delta}f(W,\delta)$ can be solved efficiently, yielding useful insights into $\Phi$. However, finding the global minimum of $\Phi$ remains NP-hard in general due to its nonconvex nature. To address this challenge, we define local surrogates for the global minimum of $\Phi$. One commonly used surrogate in nonconvex optimization is the notion of stationarity, which is suitable when $\Phi$ is differentiable. A point $W$ is an $\epsilon$-stationary point $(\epsilon\geq 0)$ of a differentiable function $\Phi$ if $\|\nabla\Phi(W)\|\leq\epsilon$.
@@
The objective function and constraint set $\left(f:\mathbb{R}^{d\times k+n}\rightarrow\mathbb{R},\Delta\subseteq\mathbb{R}^{n}\right)$ satisfy (i) $\Delta$ is a convex and bounded set with a diameter $D\geq 0$. (ii) $f$ has $\ell$-Lipchits gradients and is $\mu$-strongly concave in $\delta$. That is, for both $\in\{W,\delta\}$
$\left\|\nabla_{*}f(W,\delta)-\nabla_{*}f\left(W^{\prime},\delta^{\prime}\right)\right\|_{F}^{2}\leq\ell^{2}\left(\left\|W-W^{\prime}\right\|^{2}_{F}+\left\|\delta-\delta^{\prime}\right\|^{2}_{F}\right).$ (8)
Let $\kappa=\ell/\mu$ denote the condition number and define
@@
The following theorem characterizes the convergence rate of the proposed AdvCLIP-LoRA in Alg. 1 to find a stationary solution for $\Phi(W)$.
@@
Let Assumptions 4.1 and 4.2 hold. Moreover, assume that the low-rank matrices remain bounded by constants $c_{A}$ and $c_{B}$ in each iteration, i.e., $\|A_{t}\|_{F}\leq c_{A}$ and $\|B_{t}\|_{F}\leq c_{B}$. Then, there exists iteration $t\in\{0,\cdots,T-1\}$ for which
$\mathbb{E}\left\|\nabla\Phi\left(W_{t}\right)\right\|^{2}_{F}\leq\mathcal{O}\left(\frac{4\Delta_{\Phi}(1/\eta_{w})+\kappa\ell^{2}(c_{A}^{2}+c_{B}^{2})D^{2}}{\epsilon^{2}}\right),$ (10)
where $\eta_{w}=\Theta(\mathrm{min}\{1/\kappa\ell(c_{A}^{4}+c_{B}^{4}),1/\kappa^{2}\ell(c_{A}^{2}+c_{B}^{2}),1/(G^{2}+\kappa\ell c_{A}^{4}c_{B}^{2})^{1/2}\})$, $\eta_{\delta}=\Theta(1/\ell)$, and $\Delta_{\Phi}=\mathbb{E}\Phi(W_{0})-\mathbb{E}\Phi(W_{T+1})$. Moreover, the mini-batch size $M$ is bounded by
@@
AdvCLIP-LoRA is guaranteed to reach an $\epsilon$-stationary point of $\Phi(\cdot)$ in $\mathcal{O}(\epsilon^{-2})$ iterations, with total stochastic gradient complexity $\mathcal{O}(\epsilon^{-4})$, matching classical rates in the minimax optimization literature Lin et al. (2020).